OpenStudy (anonymous):

mr bloop is rolling steel disks down a ramp. each disk is 4cm in diameter and 5mm thick. it takes 7 seconds to go down the ramp and the distence traveled from the top to the bottom of the ramp is 119cm. how many revolutions does a disk make on the way down? round your answer to the nearest hundreath.

OpenStudy (anonymous):

Circumference= 2*pi*radius= pi*diameter= pi*4= 4pi total revolution= distance/ circumfence= 119/4pi= 9.47 revolutions